# Randomization Tests for Conditional Group Symmetry

### v1

@@ 3.1 Representatives and inversions
Let $X$ be a random vector in $\mathbb{R}^{d}$, acted on by $\text{SO}(d)$. A convenient maximal invariant is $M(X)=\|X\|$. The set of orbit representatives can be chosen to be the points on the axis with unit basis vector $\mathbf{e}_{1}=[1,0,\dotsc,0]^{\top}$. Then for each $x\in\mathbb{R}^{d}$, $\rho(x)=\|x\|\mathbf{e}_{1}$. For $d=2$, the action is free; for $d>2$, the stabilizer subgroup $\mathbf{G}_{\rho(x)}$ is the set of $d$-dimensional rotations around the axis corresponding to $\mathbf{e}_{1}$. One may construct a representative inversion function corresponding to $\rho(x)=\|x\|\mathbf{e}_{1}$ by, for example, rotating $\|x\|\mathbf{e}_{1}$ to $x$ in the 2D subspace spanned by $\nicefrac{{x}}{{\|x\|}}$ and $\mathbf{e}_{1}$. That is, let $\tilde{x}\coloneqq\nicefrac{{(x-\langle\mathbf{e}_{1},x\rangle\mathbf{e}_{1})}}{{\|x-\langle\mathbf{e}_{1},x\rangle\mathbf{e}_{1}\|}}$, so that $[\mathbf{e}_{1},\ \tilde{x}]$ is a matrix in $\mathbb{R}^{d\times 2}$ whose columns form an orthonormal basis for the 2D subspace spanned by $\nicefrac{{x}}{{\|x\|}}$ and $\mathbf{e}_{1}$. Now let $\theta_{x}$ be such that $\cos(\theta_{x})=\langle\mathbf{e}_{1},\nicefrac{{x}}{{\|x\|}}\rangle$, and $R_{\theta}$ the standard 2D rotation matrix of angle $\theta$ as defined in (5). Then the $d$-dimensional rotation defined by
$\tau(x)=\mathbf{I}_{d}-\mathbf{e}_{1}\mathbf{e}_{1}^{\top}-\tilde{x}\tilde{x}^{\top}+[\mathbf{e}_{1},\ \tilde{x}]R_{\theta_{x}}[\mathbf{e}_{1},\ \tilde{x}]^{\top}$ (10)
satisfies $\tau(x)\cdot\|x\|\mathbf{e}_{1}=x$. A sample from the corresponding inversion kernel is then generated by taking a uniform random $(d-1)$-dimensional rotation $H$ and extending it to a $d$-dimensional rotation $H^{\prime}$ that fixes $\mathbf{e}_{1}$, so that $\tau(x)H^{\prime}$ has distribution $P_{\tilde{\tau}|X}$.
@@
and so by definition, $Q$ is a maximal invariant. The subgroup that preserves the orientation of time and space (hence viewed as corresponding to physically realistic transformations) is known as the restricted Lorentz group, denoted $\mathrm{SO}^{+}(1,3)$. Mathematically, the restricted Lorentz group is a non-compact, non-commutative, connected real Lie group (a differentiable manifold) that consists of transformations $x\mapsto g_{R}g_{\beta}x$ composed of boosts $g_{\beta}$ and three-dimensional rotations $g_{R}$ that act on the directional entries $(p_{1},p_{2},p_{3})$. For $\beta\in(-1,1)$ and $R\in\textrm{SO}(3)$, transformations in $\mathrm{SO}^{+}(1,3)$ of a particle’s four-momentum $x$ are encoded by
$g_{\beta}\begin{bmatrix}E\\ p_{1}\\ p_{2}\\ p_{3}\end{bmatrix}$ $=\begin{bmatrix}(1-\beta^{2})^{-\frac{1}{2}}(E-\beta p_{1})\\ (1-\beta^{2})^{-\frac{1}{2}}(-\beta E+p_{1})\\ p_{2}\\ p_{3}\end{bmatrix}\;,$ $g_{R}x$ $=\begin{bmatrix}1&0&0&0\\ 0&R_{11}&R_{12}&R_{13}\\ 0&R_{21}&R_{22}&R_{23}\\ 0&R_{31}&R_{32}&R_{33}\end{bmatrix}x\;.$
@@
These transformations are sometimes referred to as “hyperbolic rotations” that give rise to orbits whose one-dimensional projections are hyperbolas. The set of orbit representatives can be chosen as $\rho(x)=(E_{\rho(x)},1,0,0)$, so that $[\mathbf{X}]=\{(E_{\rho},1,0,0):E_{\rho}\in\mathbb{R}_{+}\}$, with $E_{\rho(x)}^{2}-1=Q(x)$. For any $x=(E,p_{1},p_{2},p_{3})$, a sample from the inversion kernel $P_{\tilde{\tau}|X}({\,\vbox{\hbox{\tiny$\bullet$}}\,}|x)$ can be obtained in two steps: first, calculate $\beta=\frac{E_{\rho(x)}-E\sqrt{p_{1}^{2}+p_{2}^{2}+p_{3}^{2}}}{1+E^{2}}$; second, sample a 3D rotation $R$ uniformly at random from $\mathbf{G}_{(1,0,0)}$ and compute $\tau((p_{1},p_{2},p_{3}))$ as in (10). Then $x=g_{\tau((p_{1},p_{2},p_{3}))}G_{R}g_{\beta}\rho(x)$.
4 Characterizations of conditional symmetry
Definition 3.1 does not provide an actionable characterization of conditional symmetry with which to build hypothesis tests. In this section, we establish a number of characterizations of conditional symmetry. While they have immediate implications for testing, a detailed discussion on the construction of these tests will be the main focus of Section 5. The first characterization is in terms of a particular form of conditional independence that elucidates the structure of equivariant conditional distributions.
@@ 4 Characterizations of conditional symmetry
Equivariant Markov kernels arise naturally in the disintegration of jointly invariant measures. Under general conditions, a measure is jointly invariant if and only if it disintegrates into an invariant marginal measure and an equivariant Markov kernel. A measure-theoretic proof is provided by Kallenberg (2017, Theorem 7.6). Bloem-Reddy and Teh (2020) connected this to conditional independence, showing that for compact $\mathbf{G}$ and $\mathbf{G}$-invariant $P_{X}$ such that a measurable representative inversion $\tau$ exists, $P_{Y|X}$ is $\mathbf{G}$-equivariant if and only if
$X{\;\perp\!\!\!\perp\;}\tau(X)^{-1}Y\mid M(X)\;.$ (12)
@@
Informally, the relation in (12) says that given knowledge of the orbit through a maximal invariant $M(X)$, the distribution of the transformed variable $\tau(X)^{-1}Y$ does not depend on the particular value $X$ takes on its orbit. Figure 5 illustrates this intuition for a simple rotation example: the aligned conditional distributions of $Y|X=x$ for different values of $x$ on the same orbit should be the same distribution; if not, additional dependency between $Y$ and $X$ is indicated.
@@
Figure 5: Three random samples of $Y\in\mathbb{R}^{2}$ (triangle) drawn conditionally on $X\in\mathbb{R}^{2}$ (circle). If the conditional distribution $P_{Y|X}$ is $\mathrm{SO}(2)$-equivariant, then $\tau(x)_{*}P_{Y|X}({\,\vbox{\hbox{\tiny$\bullet$}}\,}|x)$ is the same for all samples.
@@
The assumptions for (12) in (Bloem-Reddy and Teh, 2020) are somewhat restrictive, as they require $\mathbf{G}$ to be compact and $P_{X}$ to be invariant. In many situations, one or both of those assumptions will not hold. The following theorem substantially generalizes (12), making it applicable to non-compact $\mathbf{G}$ and non-invariant $P_{X}$. Moreover, it does not require the existence of a representative inversion $\tau(x)$.
@@
Let $\mathbf{G}$ be a lcscH group acting on each of $\mathbf{X}$ and $\mathbf{Y}$, the action on $\mathbf{X}$ proper, with measurable inversion kernel $P_{\tilde{\tau}|X}$. The conditional distribution $P_{Y|X}$ is $\mathbf{G}$-equivariant if and only if,
$(\tilde{\tau},X){\;\perp\!\!\!\perp\;}\tilde{\tau}^{-1}Y\mid M(X)\;,\quad\text{with}\quad\tilde{\tau}\sim P_{\tilde{\tau}|X}({\,\vbox{\hbox{\tiny$\bullet$}}\,}\mid X)\;.$ (13)
@@
The proof can be found in Section B.2. We note the following special cases. If there exists a representative inversion function $\tau\colon\mathbf{X}\to\mathbf{G}$, then condition (13) reduces to (12). If the action of $\mathbf{G}$ on $\mathbf{Y}$ is trivial, then (13) reduces to $X{\;\perp\!\!\!\perp\;}Y\mid M(X)$. Finally, if the action of $\mathbf{G}$ on $\mathbf{X}$ is transitive, then (13) specializes to the (unconditional) independence.
@@
Theorem 4.1 implies that a test for the hypotheses (8) can be performed using a test for the conditional independence in (13). If the action of $\mathbf{G}$ on $\mathbf{X}$ is transitive, then $\mathbf{X}$ has a single orbit and the test simplifies to a test for (unconditional) independence. In principle, any general-purpose test for conditional independence may be used; see Li and Fan (2020) for a recent review of the literature. However, it is known that any test of size $\alpha$ over the entire null hypothesis class has no power against any alternative (Shah and Peters, 2020). The high-level solution is to restrict the null hypothesis class, which is commonly accomplished through problem-specific assumptions (Shah and Peters, 2020; Kim et al., 2022) or modeling some parts of the relevant conditional distributions (Zhang et al., 2011; Pogodin et al., 2024).
@@
One broadly useful set of tools for testing conditional independence are conditional randomization tests (Candes et al., 2018; Berrett et al., 2019; Liu et al., 2022), in which a suitable randomization of the data is used to generate a test that is similar in spirit to Algorithm 1. This motivates a refinement of the characterization of conditional symmetry in Theorem 4.1, on which we will build a conditional randomization test in the next section. The basic idea is that for a pair $(X,Y)$, the “centered” variables $\rho(X)$ and $\tilde{Y}\coloneqq\tilde{\tau}^{-1}Y$ can be used to obtain a randomized pair $(X^{\prime},Y^{\prime})\stackrel{{\scriptstyle\text{\rm\tiny d}}}{{=}}(X,Y)$. The randomization involves sampling one or both of
$\tilde{Y}\sim P_{\tilde{Y}|[X]}({\,\vbox{\hbox{\tiny$\bullet$}}\,}\mid\rho(X))\;,\quad\Gamma\sim P_{\Gamma|[X]}({\,\vbox{\hbox{\tiny$\bullet$}}\,}\mid\rho(X))\;,$ (14)
@@
with $\tilde{Y}{\;\perp\!\!\!\perp\;}\Gamma\mid\rho(X)$.
Proposition 4.2.
Let $P_{X,Y}=P_{X}\otimes P_{Y|X}$ be a probability measure on $\mathbf{X}\times\mathbf{Y}$. For a lcscH group $\mathbf{G}$ acting on $\mathbf{X}$ and $\mathbf{Y}$, with the action on $\mathbf{X}$ proper, let $P_{\Gamma|[X]}$ be the group-orbit kernel (9). Then the following statements are equivalent to $P_{Y|X}$ being $\mathbf{G}$-equivariant.
  1. E1.
    $(\rho(X),\tilde{Y},X,Y)\stackrel{{\scriptstyle\text{\rm\tiny d}}}{{=}}(\rho(X),\tilde{Y},\Gamma\rho(X),\Gamma\tilde{Y})$, with $\Gamma$ sampled as in (14).
  2. E2.
    $(\rho(X),\tilde{\tau},X,Y)\stackrel{{\scriptstyle\text{\rm\tiny d}}}{{=}}(\rho(X),\tilde{\tau},\tilde{\tau}\rho(X),\tilde{\tau}\tilde{Y})$, with $\tilde{Y}$ sampled as in (14).
  3. E3.
    $(\rho(X),X,Y)\stackrel{{\scriptstyle\text{\rm\tiny d}}}{{=}}(\rho(X),\Gamma\rho(X),\Gamma\tilde{Y})$, with $\tilde{Y},\Gamma$ sampled as in (14).
These properties follow as elaborations of Theorem 4.1, and the proof can be found in Section B.2. Taken together, they say that if we are able to sample at least one of $\Gamma$ and $\tilde{Y}$ from the correct distribution conditional on $\rho(X)$, we can test for $\mathbf{G}$-equivariance by ftesting for equality in distribution between the data and the randomization $(X^{\prime},Y^{\prime})\coloneq(\Gamma\rho(X),\Gamma\tilde{Y})$. Crucially, the randomization (and therefore the test) can be conditional on part of the observed data. The part that gets conditioned on are the variable(s) that precede $X,Y$ on the left-hand side of the distributional identity in each property. For example, Property E1 allows us to condition on $(\rho(Y),\tilde{Y})$. In each case, conditioning on part of the data restricts the null hypothesis set to agree with that part of the data, and therefore sidesteps the impossibility result of Shah and Peters (2020).
Proposition 4.2 has two special cases that are important for practical reasons. First, if $\mathbf{G}$ acts non-trivially on $\mathbf{Y}$, the data being used in the test can be reduced to $Y$ only.
Corollary 4.3.
Assume the conditions of Proposition 4.2. If $\mathbf{G}$ acts non-trivially on $\mathbf{Y}$ then the following statements are equivalent to $P_{Y|X}$ being $\mathbf{G}$-equivariant.
  1. E4.
    $(\rho(X),\tilde{Y},Y)\stackrel{{\scriptstyle\text{\rm\tiny d}}}{{=}}(\rho(X),\tilde{Y},\Gamma\tilde{Y})$, with $\Gamma$ sampled as in (14).
  2. E5.
    $(\rho(X),\tilde{\tau},Y)\stackrel{{\scriptstyle\text{\rm\tiny d}}}{{=}}(\rho(X),\tilde{\tau},\tilde{\tau}\tilde{Y})$, with $\tilde{Y}$ sampled as in (14).
  3. E6.
    $(\rho(X),Y)\stackrel{{\scriptstyle\text{\rm\tiny d}}}{{=}}(\rho(X),\Gamma\tilde{Y})$, with $\tilde{Y},\Gamma$ sampled as in (14).
Corollary 4.4.
  1. E7.
    $(\rho(X),\tilde{Y},X,Y)\stackrel{{\scriptstyle\text{\rm\tiny d}}}{{=}}(\rho(X),\tilde{Y},G\rho(X),G\tilde{Y})$ with $G\sim\lambda$.
  2. E8.
@@
Consider random vectors $X$ and $Y$ in $\mathbb{R}^{d}$, acted on by $\mathrm{SO}(d)$. As described in Example 3, a convenient maximal invariant is $M(X)=\|X\|$. Then Property E6 says that $P_{Y|X}$ is $\mathrm{SO}(d)$-equivariant if and only if $Y\stackrel{{\scriptstyle\text{\rm\tiny d}}}{{=}}\Gamma\tilde{Y}$, where $\tilde{Y}\sim P_{Y|X}({\,\vbox{\hbox{\tiny$\bullet$}}\,}\mid\|X\|\mathbf{e}_{1})$, and where $\Gamma\sim P_{\Gamma|[X]}({\,\vbox{\hbox{\tiny$\bullet$}}\,}\mid\|X\|\mathbf{e}_{1})$ is a random $d$-dimensional rotation matrix. Property E5 replaces $\Gamma$ with $\tilde{\tau}$, where $\tilde{\tau}\sim P_{\tilde{\tau}|X}({\,\vbox{\hbox{\tiny$\bullet$}}\,}|X)$ (for which the sampling procedure is described in Example 3); alternatively, Property E4 replaces $\tilde{Y}$ with $\tilde{\tau}^{-1}Y$. If $P_{Y|X}=\mathsf{N}(X,\Sigma)$, then $P_{Y|X}$ is $\mathrm{SO}(d)$-equivariant if and only if $\Sigma=\mathbf{I}_{d}$.
Example 6 (Continuation of Example 4).
The Standard Model of physics is invariant under the Lorentz group, and violations of $\mathrm{SO}^{+}(1,3)$-invariance could indicate phenomena beyond the Standard Model. (More modestly, violations of $\mathrm{SO}^{+}(1,3)$-invariance could indicate issues with high-energy physics simulators.) As an example, consider the four-momenta of particles produced by high-energy colliders such as the Large Hadron Collider (LHC). High-energy quarks produced in collisions such as those at the LHC quickly decay through a cascading process of gluon emission into collimated sprays of stable hadrons, which are subatomic particles that can be detected and measured (Salam, 2010; Bogatskiy et al., 2020). This spray is known as a jet. According to the Standard Model, the four-momenta of the constituent particles should respect the symmetries of the Lorentz group in the following sense. Let $X$ and $Y$ be the four-momenta relative to the lab frame of any two particles produced by the same collision; hence, they are stochastically dependent. The conditional distribution $P_{Y|X}$ should be $\mathrm{SO}^{+}(1,3)$-equivariant, as a Lorentz transformation of $X$ is equivalent to a change of the lab reference frame, which would then also apply to $Y$. Observe that this is not distributional invariance, because $gX\stackrel{{\scriptstyle\text{\rm\tiny d}}}{{\neq}}X$. Example 4 describes how to sample from the inversion kernel, so if samples can be generated from $X^{\prime}\sim P_{X|Q(X)}({\,\vbox{\hbox{\tiny$\bullet$}}\,}|Q(X))$, then $\Gamma=\tilde{\tau}^{\prime}$ with $\tilde{\tau}^{\prime}\sim P_{\tilde{\tau}|X}({\,\vbox{\hbox{\tiny$\bullet$}}\,}|X^{\prime})$ and $\tilde{Y}=\tilde{\tau}^{-1}Y$ with $\tilde{\tau}\sim P_{\tilde{\tau}|X}({\,\vbox{\hbox{\tiny$\bullet$}}\,}|X)$ satisfy the distributional identities above.
In the experiment in Section 7.4, we test whether this equivariance holds, using the Top Quark Tagging Reference dataset (Kasieczka et al., 2019), which consists of simulated collision events generated by PYTHIA, a Monte Carlo simulator that is widely used in high-energy physics.

### v2

@@ 3 Characterizations of conditional symmetry
Let $X$ be a random vector in $\mathbb{R}^{d}$, acted on by $\text{SO}(d)$. A convenient maximal invariant is $M(X)=\|X\|$. The set of orbit representatives can be chosen to be the points on the axis with unit basis vector $\mathbf{e}_{1}=[1,0,\dotsc,0]^{\top}$, i.e., $\rho(x)=\|x\|\mathbf{e}_{1}$ for $x\in\mathbb{R}^{d}$. For $d=2$, the action is free; for $d>2$, the stabilizer subgroup $\mathbf{G}_{\rho(x)}$ is the set of $d$-dimensional rotations around the axis corresponding to $\mathbf{e}_{1}$. One may construct a representative inversion function corresponding to $\rho(x)=\|x\|\mathbf{e}_{1}$ by, for example, rotating $\|x\|\mathbf{e}_{1}$ to $x$ in the 2D subspace spanned by $\nicefrac{{x}}{{\|x\|}}$ and $\mathbf{e}_{1}$. That is, let $\tilde{x}\coloneqq\nicefrac{{(x-\langle\mathbf{e}_{1},x\rangle\mathbf{e}_{1})}}{{\|x-\langle\mathbf{e}_{1},x\rangle\mathbf{e}_{1}\|}}$, so that $\mathbf{e}_{1},\ \tilde{x}$] is a matrix in $\mathbb{R}^{d\times 2}$ whose columns form an orthonormal basis for the 2D subspace spanned by $\nicefrac{{x}}{{\|x\|}}$ and $\mathbf{e}_{1}$. Let $\Theta_{x}$ be the 2D rotation matrix of angle $\cos^{-1}(\langle\mathbf{e}_{1},\nicefrac{{x}}{{\|x\|}}\rangle)$. Then the $d$-dimensional rotation defined by
${\psi^{\circ}}(x)=\mathbf{I}_{d}-\mathbf{e}_{1}\mathbf{e}_{1}^{\top}-\tilde{x}\tilde{x}^{\top}+[\mathbf{e}_{1},\ \tilde{x}]\Theta_{x}[\mathbf{e}_{1},\ \tilde{x}$$^{\top}$ (4)
satisfies ${\psi^{\circ}}(x)\rho(x)=x$. A sample from the corresponding inversion kernel is generated by taking a uniform random (d1)(d-1)-dimensional rotation $H$ and extending it to a $d$-dimensional rotation $H^{\prime}$ that fixes $\mathbf{e}_{1}$, so that $\Psi\stackrel{{\scriptstyle\text{\rm\tiny d}}}{{=}}{\psi^{\circ}}(x)H^{\prime}$ has distribution $P_{\Psi|X}$.
@@
and so by definition, $Q$ is a maximal invariant. The subgroup that preserves the orientation of time and space (hence viewed as corresponding to physically realistic transformations) is known as the restricted Lorentz group, denoted $\mathrm{SO}^{+}(1,3)$. The restricted Lorentz group is a non-compact, non-commutative, connected real Lie group (a differentiable manifold) that consists of transformations $x\mapsto g_{O}g_{\beta}x$ composed of boosts $g_{\beta}$ and three-dimensional rotations $g_{O}$ that act on the directional entries $(p_{1},p_{2},p_{3})$. For $\beta\in(-1,1)$ and $O\in\textrm{SO}(3)$, $\mathrm{SO}^{+}(1,3)$ transformations of a particle’s four-momentum $x$ are encoded by
$g_{\beta}\begin{bmatrix}E\\ p_{1}\\ p_{2}\\ p_{3}\end{bmatrix}$ $=\begin{bmatrix}(1-\beta^{2})^{-\frac{1}{2}}(E-\beta p_{1})\\ (1-\beta^{2})^{-\frac{1}{2}}(-\beta E+p_{1})\\ p_{2}\\ p_{3}\end{bmatrix}\;$, $g_{O}x$ $=\begin{bmatrix}1&0&0&0\\ 0&O_{11}&O_{12}&O_{13}\\ 0&O_{21}&O_{22}&O_{23}\\ 0&O_{31}&O_{32}&O_{33}\end{bmatrix}x\;.$
@@
These transformations are sometimes referred to as “hyperbolic rotations” that give rise to orbits whose one-dimensional projections are hyperbolas. The set of orbit representatives can be chosen as $\rho(x)=(E_{\rho(x)},1,0,0)$, with $E_{\rho(x)}^{2}-1=Q(x)$, so that $\mathbf{R}=\{(E,1,0,0):E\in\mathbb{R}_{+}\}$. For any $x=(E,p_{1},p_{2},p_{3})$, a sample from the inversion kernel $P_{\Psi|X}({\,\vbox{\hbox{\tiny$\bullet$}}\,}|x)$ can be obtained in two steps: first, calculate $\beta=\frac{E_{\rho(x)}-E\sqrt{p_{1}^{2}+p_{2}^{2}+p_{3}^{2}}}{1+E^{2}}$; second, sample a 3D rotation $O$ uniformly at random from $\mathbf{G}_{(1,0,0)}$ and compute ${\psi^{\circ}}((p_{1},p_{2},p_{3}))$ as in (4). Then $x=g_{{\psi^{\circ}}((p_{1},p_{2},p_{3}))}G_{O}g_{\beta}\rho(x)$.
3.1 Equivariance as conditional independence
@@
Equivariant Markov kernels arise naturally in the disintegration of jointly invariant measures. Under general conditions, a measure is jointly invariant if and only if it disintegrates into an invariant marginal measure and an equivariant Markov kernel. A measure-theoretic proof of this fact is provided by Kallenberg (2017, Theorem 7.6). Bloem-Reddy and Teh (2020) connected this to conditional independence, showing that for compact $\mathbf{G}$ acting on $\mathbf{X}$ such that a measurable representative inversion $\psi^{\circ}$} exists, and $\mathbf{G}$-invariant $P_{X}$, $P_{Y|X}$ is $\mathbf{G}$-equivariant if and only if
$X{\;\perp\!\!\!\perp\;}{\psi^{\circ}}(X)^{-1}Y\mid M(X)\;.$ (5)
@@
Informally, the relation in (5) says that given knowledge of the orbit through a maximal invariant $M(X)$, the distribution of the transformed variable ${\psi^{\circ}}(X)^{-1}Y$ does not depend on the particular value $X$ takes on its orbit. Figure 3 illustrates this intuition for the simple example of 2D rotations: the aligned conditional distributions of $Y|X=x$ for different values of $x$ on the same orbit should be the same. If they are not, additional dependence between $Y$ and $X$ is indicated.
@@
Figure 3: Three random samples of $Y\in\mathbb{R}^{2}$ (triangle) drawn conditionally on $X\in\mathbb{R}^{2}$ (circle). If the conditional distribution $P_{Y|X}$ is $\mathrm{SO}(2)$-equivariant, then $P_{Y|X}({\psi^{\circ}}(x)^{-1}{\,\vbox{\hbox{\tiny$\bullet$}}\,}|x)$ is the same for all samples.
@@
The assumptions for (5) in (Bloem-Reddy and Teh, 2020) are somewhat restrictive, as they require $\mathbf{G}$ to be compact and $P_{X}$ to be invariant. In many situations, one or both of those assumptions will not hold. The following theorem substantially generalizes (5), making it applicable to non-compact $\mathbf{G}$ and non-invariant $P_{X}$. Moreover, it does not require the existence of a representative inversion ${\psi^{\circ}}(x)$, so that the action of $\mathbf{G}$ on $\mathbf{X}$ may not be free.
@@
Let $\mathbf{G}$ be a lcscH group acting on each of $\mathbf{X}$ and $\mathbf{Y}$, the action on $\mathbf{X}$ proper, with measurable inversion kernel $P_{\Psi|X}$. The conditional distribution $P_{Y|X}$ is $\mathbf{G}$-equivariant if and only if,
$(\Psi,X){\;\perp\!\!\!\perp\;}\Psi^{-1}Y\mid M(X)\;,\quad\text{with}\quad\Psi\mid X\sim P_{\Psi|X}\;.$ (6)
@@
The proof can be found in Section C.1. We note the following special cases. If there exists a representative inversion function ${\psi^{\circ}}\colon\mathbf{X}\to\mathbf{G}$, then condition (6) reduces to (5). If the action of $\mathbf{G}$ on $\mathbf{Y}$ is trivial, then (6) reduces to $X{\;\perp\!\!\!\perp\;}Y\mid M(X)$. Finally, if the action of $\mathbf{G}$ on $\mathbf{X}$ is transitive, as in Example 1, then (6) specializes to the (unconditional) independence.
@@
Theorem 3.1 implies that a test for the hypotheses (3) can be performed using a test for the conditional independence in (6). If the action of $\mathbf{G}$ on $\mathbf{X}$ is transitive, then $\mathbf{X}$ has a single orbit and the test simplifies to a test for (unconditional) independence. In principle, any general-purpose test for conditional independence may be used. However, it is known that any test of level $\alpha$ over the entire null hypothesis class has no power against any alternative (Shah and Peters, 2020). The high-level solution is to restrict the null hypothesis class, which is commonly accomplished through problem-specific assumptions (Shah and Peters, 2020; Kim et al., 2022) or modeling some parts of the relevant conditional distributions (Zhang et al., 2011; Pogodin et al., 2024).
3.2 Distributional identities for conditional randomization
@@
One broadly useful set of tools for testing conditional independence are conditional randomization tests (Candès et al., 2018; Berrett et al., 2019; Liu et al., 2022), in which a suitable randomization of the data is used to generate a test that is similar in spirit to Algorithm 1. This motivates a refinement of the characterization of conditional symmetry in Theorem 3.1, on which we will build a conditional randomization test in the next section. The basic idea is that for a pair $(X,Y)$, the “centered” variables $\rho(X)$ and $\tilde{Y}\coloneqq\Psi^{-1}Y$, with $\Psi|X\sim P_{\Psi|X}$, can be used to obtain a randomized pair $(X^{\prime},Y^{\prime})\stackrel{{\scriptstyle\text{\rm\tiny d}}}{{=}}(X,Y)$. The randomization involves sampling one or both of
$\tilde{Y}\mid\rho(X)\sim P_{\tilde{Y}|R}\;,\quad\Gamma\mid\rho(X)\sim P_{\Gamma|R}\;,$ (7)
@@
with $\tilde{Y}{\;\perp\!\!\!\perp\;}\Gamma\mid\rho(X)$. Here, $\Gamma$ denotes the random variable with conditional distribution
$P_{\Gamma|R}({\,\vbox{\hbox{\tiny$\bullet$}}\,}\mid r)=\int_{\mathbf{X}}P_{\Psi|X}({\,\vbox{\hbox{\tiny$\bullet$}}\,}\mid x^{\prime})P_{X|R}(dx^{\prime}\mid r)\;,$ (8)
where $P_{X|R}$ is the conditional distribution of $X$ given an orbit, represented by the random representative $R$ (or more generally, any maximal invariant $M$). We refer to $P_{\Gamma|R}$ as the group-orbit kernel as it is the distribution on $\mathbf{G}$ induced by restricting $P_{X}$ to the orbit of $R$. The group-orbit kernel has the property that if $\Gamma|\rho(X)\sim P_{\Gamma|R}$, then $(\rho(X),X)\stackrel{{\scriptstyle\text{\rm\tiny d}}}{{=}}(\rho(X),\Gamma\rho(X))$. (See Lemma A.2 for a proof.)
In Section C.2, we prove a number of distributional identities involving these randomizations and that underpin the tests we construct in Section 4. In particular, with $\Gamma$ sampled as in (7), $P_{Y|X}$ is $\mathbf{G}$-equivariant if and only if
[…]
which implies that as long as $\mathbf{G}$ acts non-trivially on $\mathbf{Y}$, the distributional comparison can be carried out using only observed $Y$ and randomized $Y^{\prime}=\Gamma\tilde{Y}$.
@@
Example 4.
Consider random vectors $X$ and $Y$ in $\mathbb{R}^{d}$, acted on by $\mathrm{SO}(d)$. As described in Example 2, a convenient maximal invariant is $M(X)=\|X\|$, with $\rho(X)=\|X\|\mathbf{e}_{1}$. Then (10) says that $P_{Y|X}$ is $\mathrm{SO}(d)$-equivariant if and only if $Y\stackrel{{\scriptstyle\text{\rm\tiny d}}}{{=}}\Gamma\tilde{Y}$, where $\Gamma|\rho(X)\sim P_{\Gamma|R}$ is a random $d$-dimensional rotation matrix that satisfies $\Gamma\rho(X)\stackrel{{\scriptstyle\text{\rm\tiny d}}}{{=}}X$.
@@
The Standard Model of physics is invariant under the Lorentz group, and violations of $\mathrm{SO}^{+}(1,3)$-invariance could indicate phenomena beyond the Standard Model. (More modestly, violations of $\mathrm{SO}^{+}(1,3)$-invariance could indicate issues with high-energy physics simulators.) As an example, consider the four-momenta of particles produced by high-energy colliders such as the Large Hadron Collider (LHC). High-energy quarks produced in collisions such as those at the LHC quickly decay through a cascading process of gluon emission into collimated sprays of stable hadrons, which are subatomic particles that can be detected and measured (Salam, 2010; Bogatskiy et al., 2020). This spray is known as a jet. According to the Standard Model, the four-momenta of the constituent particles should respect the symmetries of the Lorentz group in the following sense. Let $X$ and $Y$ be the four-momenta relative to the lab frame of any two particles produced by the same collision; hence, they are stochastically dependent. The conditional distribution $P_{Y|X}$ should be $\mathrm{SO}^{+}(1,3)$-equivariant, as a Lorentz transformation of $X$ is equivalent to a change of the lab reference frame, which would then also apply to $Y$. Observe that this is not distributional invariance, because $gX\stackrel{{\scriptstyle\text{\rm\tiny d}}}{{\neq}}X$. Example 3 describes how to sample from the inversion kernel, so if samples can be generated from $X^{\prime}|\rho(X)\sim P_{X|R}$ (with $\rho$ as described in Example 3), then $\Gamma=\Psi^{\prime}$ with $\Psi^{\prime}|X^{\prime}\sim P_{\Psi|X}$ and $\tilde{Y}=\Psi^{-1}Y$ with $\Psi|X\sim P_{\Psi|X}$ satisfy the distributional identities above.